A new cruise ship line has just launched 3 new​ ships: the Pacific​ Paradise, the Caribbean​ Paradise, and the Mediterranean Paradise. The Caribbean Paradise has 36 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 37 fewer deluxe staterooms than four times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 605.
Answer by Boreal(15235)

P=x
C=x+36
M=4x-37
They add to 605
6x-1
605
6x=606
x=101 Pacific
x+36=137 Caribbean
4x-37=367 Mediterranean